Help Save & Ex Nations Capital Fitness, Inc. operates a chain of fitness centers In the Washington, D.Cc, area. The firms controller is accumulating data to be used in preparing its annual profit plan for the coming year. The cost behavior pattern of the firms equipment maintenance costs must be determined. The accounting staff has suggested the use of the high-low method to develop an equation, In the form of Y -a + bx, for maintenance costs. Data regarding the maintenance hours and costs last year are as follows: the maintenance hours and costs for MaintenanceMaintenance Costs January February March April May June July August Septembe 370 October November December340 30 470 260 480 350 430 330 440 490 5,045 4,250 2,750 4,320 3,020 4,160 3,080 3,590 3,970 3,280 3,200 3,080 $43,745 370 4,860 Total Average 405* Rounded 4-a. Compute the variable cost per hour and the fixed cost per hour at 610 hours of activity. (Round your answers to 2 decimal places Variable cost per hour Fixed cost per hour < Prey 9 of 14 İİİ Next >
4-b. Why is the fixed cost per hour possibly misleading? Because it will change as the number of hours changes. O Because it will not change as the number of hours changes. K Prev 9 of 1
Answer #1

Variable and fixed costs @ 610 hours of capacity

Variable cost per hour: $8.5

Fixed cost per hour:$0.88 ($540/610 hours)

We have,
at highest activity: x2 = 530; y2 = $5045
at lowest activity: x1 = 260; y1 = $2750

Variable Cost per Unit = ($5045 − $2750) ÷ (530 − 260) = $8.5 per unit
Total Fixed Cost = $5045 − ($8.5 × 530) = $2750 − ($8.5 × 260) = $540
Cost Volume Formula: y = $540 + 8.5x

4-b) Because it will change as a number of hours change

Fixed cost remain fixed in value and may change per hour cost  due to number of hours increased